Two metal spheres, each of mass 10 g and initially at rest, are dropped from a height of 10 m in an evacuated chamber. One sphere has a charge of +100 μC, and the other has a charge –100 μC. Find the difference in final speeds of the two spheres (the speeds just before each one hits the ground). Assume that there is a downward electric field of about 150 V/m near the Earth’s surface.
Answer #1

acceleration of negative charge

a- = g + qE/m

a- = 9.8 + 100* 10^-6* 150 / 0.01

a- = 11.3 m/s^2

velocity reaching the ground

v-^2 = 2 (a-) d = 2* 11.3* 10

v- = 15.03 m/s

acceleration of positive charge

a+ = 9.8 - 100* 10^-6* 150 / 0.01

a+ = 8.3 m/s^2

velocity reaching the ground

v+^2 = 2 * 8.3* 10

v+ = 12.884 m/s

difference in velocity

v = 2.146 m/s

=======

Comment before rate in case any doubt, will reply for sure.. 5